# Cristiá, Capozucca and Rossi

### v1

@@ 1 Introduction
The work presented in this paper is aligned with the following quotations (Garavel et al.,, 2020).
@@
That is, this paper presents $\{log\}$ (read ‘setlog’) (Rossi and Cristiá,, 2000; Dovier et al.,, 2000), a programming language that has the look-and-feel of a formal specification language which in turn is supported by a proof system. The core of $\{log\}$ is a Constraint Logic Programming (CLP) language and interpreter written in SWI-Prolog implementing a constraint satisfiability solver for set theory. The satisfiability solver is a rewriting system implementing decision procedures for several fragments of set theory (e.g., integer intervals (Cristiá and Rossi, 2024c, ) and restricted quantifiers (Cristiá and Rossi, 2024d, )). In this way, $\{log\}$ implements declarative programming in the form of set programming Schwartz et al., (1986); Cantone et al., (2001). As such $\{log\}$ has been applied to several case studies showing that the satisfiability solver works in practice (Cristiá and Rossi, 2021a, ; Cristiá and Rossi, 2021c, ; Cristiá and Rossi, 2024a, ; Cristiá et al.,, 2023; Capozucca et al.,, 2024).
@@
For these reasons, over the years, $\{log\}$ has been extended with some components (also implemented in SWI-Prolog) that constrain programmers to write code that is easier to verify. This programming style within $\{log\}$ defines a language for the description of state machines in terms of first-order logic (FOL), set theory and integer arithmetic—much as in formal notations such as B (Abrial,, 1996) and Z (Woodcock and Davies,, 1996). In this way: (i) the state machine language allows for the automatic generation of verification conditions (VC); (ii) these VC are automatically fed into the satisfiability solver which informs the user if they could be discharged or not (in which case, a counterexample is generated); (iii) state machines can still be executed as regular $\{log\}$ code thus allowing users to use them as working functional prototypes before attempting any serious proof campaign; and (iv) users can generate test cases to test a future, more efficient implementation written in some other programming language. This paper makes a comprehensive presentation of these extensions aiming at showing how theoretical contributions in the fields of CLP and set programming can be turned into a feasible tool aimed at formal verification.
As the above quotations suggest, there are other systems where programming and proof coexist. Four conspicuous, widely known, successful representatives of this kind are Agda (Logic and Types group,, 2004), Dafny (Leino,, 2010), F* (read ‘f-star’) (Microsoft Research and INRIA,, 2015) and Why3 (Toccata group,, 2012). All of them have been applied to a range of problems, some times of industrial size and complexity. Agda is a typed functional programming language and proof assistant based on intuitionistic type theory. The functional language provides usual functional features such as inductive types, pattern matching, and so on. Proofs are interactive and written in a functional programming style that, unlike Coq (now renamed as Rocq) (Bertot and Castéran,, 2004), does not provide proof tactics. Programmers can, nonetheless, implement Agda functions that return a proof of some property [^1] Writing such proofs requires an unusual degree of training in formal software development.. These functions embody sort of proof tactics. Later, these functions are run during type checking: if one of them fails, then type checking fails. This means that the program does not verify a certain property. As a form of proof automation Agda provides proof search by enumerating possible proof terms that may fit in the specification. Dafny combines ideas from the functional and imperative paradigms, including some support for object-oriented programming. Programs are annotated with specifications through preconditions, postconditions, loop invariants, etc. The specification language has the look-and-feel of a programming language including mathematical integers, bit-vectors, sequences, sets, induction, and so on. Proof obligations are automatically discharged, given sufficient specification, by the Z3 SMT solver (Microsoft Research,, 2008). Sufficient specification means, in general, to write sufficient proof steps (e.g., the inductive step in an inductive proof) the remainder of which are discharged automatically. F* supports purely functional programming as well as effectful programming (McBride and Paterson,, 2008). Its powerful (dependent) type system allows to describe precise program specifications including security and functional properties. The typechecker will check whether or not the program verifies its specification by discharging proofs interactively written by the programmer (using tactics, metaprogramming and symbolic computation) or by calling the Z3 SMT solver. F* shares some features with Agda. Why3 is a platform for deductive program verification. It provides WhyML, a specification and programming language. Why3 uses many external theorem provers (e.g., Alt-Ergo (OCamlPro,, 2006), CVC4-5 (Barrett et al.,, 2011; Barbosa et al.,, 2022) and E-prover (Schulz,, 1999; Schulz et al.,, 2019)) to discharge proof obligations. The system includes a library of mathematical theories (e.g., arithmetic and sets) and some programming data structures (e.g., arrays and hash tables).
@@
Having a different origin than the systems mentioned above and not being strictly programming and proof systems, we should also mention Atelier-B (Clearsy,, 2009) and Rodin (Dep. Sys. and Soft. Eng. Grp.,, 2004). These tools are meant to be used to write and verify B and Event-B (Abrial,, 2010) specifications, respectively. In both languages, specifications take the form of state machines described by means of FOL and set theory. This is the main point of contact with $\{log\}$. Both tools rely on automatic and interactive provers, including internal ones and external SMT solvers such as VeriT (Déharbe and Fontaine,, 2009) and CVC4. B and Event-B specifications can be executed (or animated) if the ProB tool (Leuschel,, 2003) is available in Atelier-B and Rodin. ProB is a constraint solver developed in Prolog. Unlike $\{log\}$, in general, ProB cannot prove properties although it can disprove them.
@@
None of these extensions have been published before nor they have been presented as an integrated verification framework with $\{log\}$ at its center.
@@ 2 The $\{log\}$ CLP language and satisfiability solver
$\{log\}$ is a publicly available constraint satisfiability solver and a declarative set-based, constraint-based programming language implemented in SWI-Prolog (Rossi and Cristiá,, 2000; Dovier et al.,, 2000). $\{log\}$ is deeply rooted in the work on Computable Set Theory (Cantone et al.,, 1989), combined with the ideas put forward by the set-based programming language SETL (Schwartz et al.,, 1986).
$\{log\}$ implements various decision procedures for different theories on the domain of finite sets and integer numbers. Specifically, $\{log\}$ implements decision procedures for: (1) The theory of hereditarily finite sets, i.e., finitely nested sets that are finite at each level of nesting (Dovier et al.,, 2000); (2) A very expressive fragment of the theory of finite set relation algebras (Cristiá and Rossi,, 2020, 2018); (3) Theory (1) extended with restricted intensional sets (Cristiá and Rossi, 2021b, ); (4) Theory (1) extended with cardinality constraints (Cristiá and Rossi,, 2023); (5) Theory (4) extended with integer intervals (Cristiá and Rossi, 2024c, ); (6) Quantifier-free, decidable languages extended with restricted quantifiers (Cristiá and Rossi, 2024d, ); and (7) The theory of linear integer arithmetic, by integrating an existing decision procedure for this theory. All these procedures constitute the core of the $\{log\}$ tool. Several in-depth empirical evaluations provide evidence that $\{log\}$ is able to solve non-trivial problems (Cristiá and Rossi,, 2020, 2018; Cristiá and Rossi, 2021b, ; Cristiá et al.,, 2013); in particular as an automated verifier of security properties (Cristiá and Rossi, 2021a, ; Cristiá and Rossi, 2021c, ; Cristiá et al.,, 2023; Capozucca et al.,, 2024).
@@
Below we introduce $\{log\}$. The presentation is brief and user-oriented. A complete, user-oriented presentation can be found in the $\{log\}$ user’s manual (Rossi and Cristiá,, 2025). A formal presentation can be found in the papers cited in the previous paragraph, including formal syntax and semantics; and soundness, completeness and termination proofs. We assume a basic knowledge of (constraint) logic programming and Prolog.
@@
The $\{log\}$ implementation of $smin$ is the following predicate:
smin(S,M) :- M in S & foreach(X in S, M =< X). (2)Right at first glance the differences between specification and program look minimal. For instance, we write M instead of $m$ to denote a variable (as in Prolog), in denotes set membership ($\in$), & denotes conjunction ($\land$), and foreach denotes the restricted universal quantifier ($\forall x\in S:\dots$). Indeed, the following is another implementation for $smin$:
@@
1{log}=> smin(S,10).2  S = {10 / _N1}3  Constraint: set(_N1), foreach(_X1 in _N1, 10 =< _X1)4{log}=> groundsol.5{log}=> smin(S,10).6  S = {10}In the first run (line 1), $\{log\}$ states that for 10 to be the minimum of S, then S must be a set of the form {10/_N1} meaning that 10 belongs to S and that all of its other possible elements (represented by set _N1) are grater than or equal to 10. Again, the constraint is always satisfiable—in this case by substituting _N1 by the empty set, which is the solution computed after the activation of groundsol (lines 4-6). The term {10/_N1} corresponds to $\{10\}\cup\_N1$ what implies that {10/{}} equals {10}.
@@
Hence, we can use smin as both a formula or specification and as a program or implementation. The same piece of code, no need for text describing the specification and another text describing the implementation, no need for program annotations. We call this the program-formula duality or the implementation-specification duality. Some times we refer to $\{log\}$ code as forgrams, a portmanteau word resulting from the combination of formula and program. Therefore, when $\{log\}$ programmers write code they are writing a program as a formula (Apt and Bezem,, 1999). The same solver executes a program and proves properties true of it, no need for external solvers or theorem provers. Actually, $\{log\}$ cannot be divided into a component running programs and a component proving properties.
Although some of these ideas are not new to the CLP community, it is our understanding that no other tool features set programming based on decision procedures implemented with CLP properties—e.g. the program-formula duality.
@@ 2.1 Extensional sets, relations and functions
[…]
@@
The constraint size(A,N) denotes $|A|=N$, i.e., set cardinality. The first argument can be an extensional set or the empty set; the second one can only be a variable or an integer number, but it can participate in integer constraints. Hence, $\{log\}$ can compute the solutions (or lack thereof) of444is is the standard Prolog predicate for the evaluation of integer expressions.:
@@
Functions are a subclass of binary relations—i.e., functions are sets of ordered pairs. The constraint pfun(F) states that the binary relation F is a function. $\{log\}$ provides two notions of function application: if pfun(F) holds then apply(F,X,Y) is satisfied iff $X\in dom(F)\land F(X)=Y$; in turn, applyTo(F,X,Y) denotes $\{(X,X)\};F=\{(X,Y)\}$, for any binary relation $F$.
@@
2.2 More set terms
Two more set terms are available in $\{log\}$. Terms of the form int(m,n), with both arguments either variables or integer numbers, denote integer intervals—i.e., sets of the form $[m,n]\cap\mathbb{Z}$. Interval limits can participate in integer constraints. Intervals can be arguments to all the classic set operators including cardinality—but in general they cannot be passed as arguments to relational operators. In this way $\{log\}$ is able to compute solutions (or lack thereof) for goals such as:
@@ 2.3 Negation
Negation in $\{log\}$ is a delicate matter as it is in logic programming in general (Apt and Bol,, 1994). The problem with negating a $\{log\}$ formula is that its result may not be a $\{log\}$ formula. Consider the following alternative implementation of $smin$ (1):
@@
were $X$ is an existential variable. The point is that, in general, neg(smin2) cannot be automatically computed. Fortunately, in the case of $smin$, since its encoding with smin in (1) contains no existential variables, $\{log\}$ is able to automatically compute neg(smin):
@@
property(R,S) :- dom(R,Dr) & dom(S,Ds) & subset(Dr,Ds).where dom(R,D) encodes the domain of relation R, i.e., $dom(R)=D$. Clearly, property introduces two existential variables (Dr and Ds) thus making it impossible to use neg to compute its negation. However, the nature of these variables is different from Max in the $smin$ example. Here, these variables define the domain of R and S. In other words, Dr and Ds are sort of names for the expressions $dom(R)$ and $dom(S)$. This means that the negation of property is not:
@@
It makes no sense to state $dom(R)\neq Dr$ for all $Dr$ because it would mean that the domain of $R$ does not exist. For cases like this $\{log\}$ provides the Let predicate (Cristiá and Rossi, 2024d, , Sect. 5):
@@
$\equiv\phi[e_{1}/x_{1},\dots,e_{n}/x_{n}]$ [by one-point rule (Woodcock and Davies,, 1996, Sect 4.2)]
@@
[…]

### v2

@@ 1 Introduction
The work presented in this paper is aligned with the following quotations (Garavel et al., 2020).
@@
Xavier Leroy
That is, this paper presents $\{log\}$ (read ‘setlog’) (Rossi and Cristiá, 2000; Dovier et al., 2000), a programming language that has the look-and-feel of a formal specification language which in turn is supported by a proof system. The core of $\{log\}$ is a Constraint Logic Programming (CLP) language and interpreter written in SWI-Prolog implementing a constraint satisfiability solver for set theory. The satisfiability solver is a rewriting system implementing decision procedures for several fragments of set theory (e.g., integer intervals (Cristiá and Rossi, 2024c ) and restricted quantifiers (Cristiá and Rossi, 2024d )). In this way, $\{log\}$ implements declarative programming in the form of set programming Schwartz et al., (1986); Cantone et al., (2001). As such $\{log\}$ has been applied to several case studies showing that the satisfiability solver works in practice (Cristiá and Rossi, 2021a ; Cristiá and Rossi, 2021c ; Cristiá and Rossi, 2024a ; Cristiá et al., 2023; Capozucca et al., 2024).
@@
For these reasons, over the years, $\{log\}$ has been extended with some components (also implemented in SWI-Prolog) that constrain programmers to write code that is easier to verify. This programming style within $\{log\}$ defines a language for the description of state machines in terms of first-order logic (FOL), set theory and integer arithmetic—much as in formal notations such as B (Abrial, 1996) and Z (Woodcock and Davies, 1996). In this way: (i) the state machine language allows for the automatic generation of verification conditions (VC); (ii) these VC are automatically fed into the satisfiability solver which informs the user if they could be discharged or not (in which case, a counterexample is generated); (iii) state machines can still be executed as regular $\{log\}$ code thus allowing users to use them as working functional prototypes before attempting any serious proof campaign; and (iv) users can generate test cases to test a future, more efficient implementation written in some other programming language. This paper makes a comprehensive presentation of these extensions aiming at showing how theoretical contributions in the fields of CLP and set programming can be turned into a feasible tool aimed at formal verification.
As the above quotations suggest, there are other systems where programming and proof coexist. Four conspicuous, widely known, successful representatives of this kind are Agda (Norell, 2008; Logic and Types group, 2004), Dafny (Leino, 2010a ; Leino, 2010b ), F* (read ‘f-star’) (Swamy et al., 2016; Microsoft Research and INRIA, 2015) and Why3 (Bobot et al., 2011; Toccata group, 2012). All of them have been applied to a range of problems, some times of industrial size and complexity.
Agda is a typed functional programming language and proof assistant based on intuitionistic type theory. The functional language provides usual functional features such as inductive types, pattern matching, and so on. Proofs are interactive and written in a functional programming style that, unlike Coq (now renamed as Rocq) (Bertot and Castéran, 2004), does not provide proof tactics. Programmers can, nonetheless, implement Agda functions that return a proof of some property [^1] Writing such proofs requires an unusual degree of training in formal software development.. These functions embody sort of proof tactics. Later, these functions are run during type checking: if one of them fails, then type checking fails. This means that the program does not verify a certain property. As a form of proof automation Agda provides proof search by enumerating possible proof terms that may fit in the specification.
Dafny combines ideas from the functional and imperative paradigms, including some support for object-oriented programming. Programs are annotated with specifications through preconditions, postconditions, loop invariants, etc. The specification language has the look-and-feel of a programming language including mathematical integers, bit-vectors, sequences, sets, induction, and so on. Proof obligations are automatically discharged, given sufficient specification, by the Z3 SMT solver (de Moura and Bjørner, 2008; Microsoft Research, 2008). Sufficient specification means, in general, to write sufficient proof steps (e.g., the inductive step in an inductive proof) the remainder of which are discharged automatically.
F* supports purely functional programming as well as effectful programming (McBride and Paterson, 2008). Its powerful (dependent) type system allows to describe precise program specifications including security and functional properties. The typechecker will check whether or not the program verifies its specification by discharging proofs interactively written by the programmer (using tactics, metaprogramming and symbolic computation) or by calling the Z3 SMT solver.
Why3 is a platform for deductive program verification. It provides WhyML, a specification and programming language. Why3 uses many external theorem provers (e.g., Alt-Ergo (Conchon et al., 2013; OCamlPro, 2006), CVC4-5 (Barrett et al., 2011; Barbosa et al., 2022) and E-prover (Schulz, 1999; Schulz et al., 2019)) to discharge proof obligations. The system includes a library of mathematical theories (e.g., arithmetic and sets) and some programming data structures (e.g., arrays and hash tables).
@@
Having a different origin than the systems mentioned above and not being strictly programming and proof systems, we should also mention Atelier-B (Lecomte et al., 2016; Clearsy, 2009), Rodin (Abrial et al., 2010; Dep. Sys. and Soft. Eng. Grp., 2004) and Alloy (Jackson, 2019). Atelier-B and Rodin are meant to be used to write and verify B and Event-B (Abrial, 2010) specifications. In both languages, specifications take the form of state machines described by means of FOL and set theory. This is the main point of contact with $\{log\}$. Both tools rely on automatic and interactive provers, including internal ones and external SMT solvers such as VeriT (Bouton et al., 2009; Déharbe and Fontaine, 2009) and CVC4. B and Event-B specifications can be executed (or animated) if the ProB tool (Leuschel and Butler, 2008; Leuschel, 2003) is available in Atelier-B and Rodin. ProB is a constraint solver developed in Prolog. Unlike $\{log\}$, in general, ProB cannot prove properties although it can disprove them.
While earlier versions of Alloy focused mainly on describing and analyzing software structures using FOL and set theory, Alloy 6 also natively supports behavioral modeling. It provides a unified language for both specifying the system and its expected properties—via linear temporal logic. A key difference is that Alloy 6 achieves completeness with respect to trace size, but remains bounded in relation to the cardinality of sets, a limitation that does not apply to $\{log\}$.
@@
None of these extensions have been published before nor presented as an integrated verification framework with $\{log\}$ at its center. Although state machines can be specified in $\{log\}$ without the language presented here (Cristiá and Rossi, 2021a ; Cristiá and Rossi, 2021c ; Cristiá et al., 2023), in that case, verification conditions cannot be automatically generated and the Next environment and automated test case generation are harder to implement, if possible. Hence, the state machine language is the enabler for the other features listed above. All the new features are implemented with +3 KLOC or +129 Kb of SWI-Prolog code distributed across two main files (setlog_vcg.pl and setlog_ttf.pl). [^2] The $\{log\}$ distribution comprises some other files making the core of the solver.
@@ 2 The $\{log\}$ CLP language and satisfiability solver
$\{log\}$ is a publicly available constraint satisfiability solver and a declarative set-based, constraint-based programming language implemented in SWI-Prolog (Rossi and Cristiá, 2000; Dovier et al., 2000). $\{log\}$ is deeply rooted in the work on Computable Set Theory (Cantone et al., 1989), combined with the ideas put forward by the set-based programming language SETL (Schwartz et al., 1986).
$\{log\}$ implements various decision procedures for different theories on the domain of finite sets and integer numbers. Specifically, $\{log\}$ implements decision procedures for: (1) The theory of hereditarily finite sets, i.e., finitely nested sets that are finite at each level of nesting (Dovier et al., 2000); (2) A very expressive fragment of the theory of finite set relation algebras (Cristiá and Rossi, 2020; 2018); (3) Theory (1) extended with restricted intensional sets (Cristiá and Rossi, 2021b ); (4) Theory (1) extended with cardinality constraints (Cristiá and Rossi, 2023); (5) Theory (4) extended with integer intervals (Cristiá and Rossi, 2024c ); (6) Quantifier-free, decidable languages extended with restricted quantifiers (Cristiá and Rossi, 2024d ); and (7) The theory of linear integer arithmetic, by integrating an existing decision procedure for this theory. All these procedures constitute the core of the $\{log\}$ tool. Several in-depth empirical evaluations provide evidence that $\{log\}$ is able to solve non-trivial problems (Cristiá and Rossi, 2020; 2018; Cristiá and Rossi, 2021b ; Cristiá et al., 2013); in particular as an automated verifier of security properties (Cristiá and Rossi, 2021a ; Cristiá and Rossi, 2021c ; Cristiá et al., 2023; Capozucca et al., 2024).
@@
Below we introduce $\{log\}$. The presentation is brief and user-oriented. A complete, user-oriented presentation can be found in the $\{log\}$ user’s manual (Rossi and Cristiá, 2025). A formal presentation can be found in the papers cited in the previous paragraph, including formal syntax and semantics; and soundness, completeness and termination proofs. We assume a basic knowledge of (constraint) logic programming and Prolog.
@@
The $\{log\}$ implementation of sminsmin is the following predicate:
   smin(S,M) :- M in S & foreach(X in S, M =< X). (2)Right at first glance the differences between specification and program look minimal. For instance, we write M instead of $m$ to denote a variable (as in Prolog), in denotes set membership ($\in$), & denotes conjunction ($\land$), and foreach denotes the restricted universal quantifier ($\forall x\in S:\dots$). Indeed, the following is another implementation for sminsmin:
@@
   1  {log}=> smin(S,10).   2   S = {10 / _N1}   3   Constraint: set(_N1), foreach(_X1 in _N1, 10 =< _X1)   4  {log}=> groundsol.   5  {log}=> smin(S,10).   6   S = {10}In the first run (line 1), $\{log\}$ states that for 10 to be the minimum of S, then S must be a set of the form {10/_N1} meaning that 10 belongs to S and that all of its other possible elements (represented by set _N1) are greater than or equal to 10. Again, the constraint is satisfiable—in this case by substituting _N1 by the empty set, which is the solution computed after the activation of groundsol (lines 4-6). The term {10/_N1} corresponds to $\{10\}\cup\_N1$ what implies that {10/{}} equals {10}.
@@
Hence, we can use smin as both a formula or specification and as a program or implementation. The same piece of code, no need for text describing the specification and another text describing the implementation, no need for program annotations. We call this the program-formula duality or the implementation-specification duality. Some times we refer to $\{log\}$ code as forgrams, a portmanteau word resulting from the combination of formula and program. Therefore, when $\{log\}$ programmers write code they are writing a program as a formula (Apt and Bezem, 1999). The same solver executes a program and proves properties true of it, no need for external solvers or theorem provers, at least for its decidable fragments. Actually, $\{log\}$ cannot be divided into a component running programs and a component proving properties.
Although some of these ideas are not new to the CLP community, it is our understanding that no other tool features set programming based on decision procedures implemented with CLP properties—e.g., the program-formula duality.
@@ 2.1 Extensional sets, relations and functions
The term $\{x/A\}$ is called extensional set; the empty set is denoted by $\{\}$. These are the basic constructors for sets; they can be arguments to all classic set operators—union, intersection, cardinality, etc. Set operators are provided as constraints. For example, un(A,B,C) denotes $C=A\cup B$, inters(A,B,C) denotes $C=A\cap B$, and diff(A,B,C) denotes $C=A\setminus B$. Then, we can compute the union of two sets even when they are partially specified [^4] For brevity, only the first solution is shown.:
@@
The constraint size(A,N) denotes $|A|=N$, i.e., set cardinality. The first argument can be an extensional set or the empty set; the second one can only be a variable or an integer number, but it can participate in integer constraints. Hence, $\{log\}$ can compute the solutions (or lack thereof) of555isis the standard Prolog predicate for the evaluation of integer expressions.:
@@
Functions are a subclass of binary relations—i.e., functions are sets of ordered pairs. The constraint pfun(F) states that the binary relation F is a function. $\{log\}$ provides two notions of function application: if pfun(F) holds then apply(F,X,Y) is satisfied iff $X\in dom(F)\land F(X)=Y$; in turn, applyTo(F,X,Y) denotes $\{(X,X)\};F=\{(X,Y)\}$, for any binary relation $F$. That is, applyTo(F,X,Y) is true if F is a binary relation containing exactly one pair whose first component is X and whose second component is Y—meaning that F is not necessarily a function.
@@
2.2 Integer intervals and restricted intensional sets
Two more set terms are available in $\{log\}$. Terms of the form int(m,n), with both arguments either variables or integer numbers, denote integer intervals—i.e., sets of the form $m,n]\cap\mathbb{Z}$. Interval limits can participate in integer constraints. Intervals can be arguments to all the classic set operators including cardinality—but in general they cannot be passed as arguments to relational operators. In this way $\{log\}$ is able to compute solutions (or lack thereof) for goals such as:
@@ 2.3 Negation
Negation in $\{log\}$ is a delicate matter as it is in logic programming in general (Apt and Bol, 1994). The problem with negating a $\{log\}$ formula is that its result may not be a $\{log\}$ formula. Consider the following alternative implementation of sminsmin (1):
@@
were $X$ is an existential variable. The point is that, in general, neg(smin2) cannot be automatically computed. Fortunately, in the case of sminsmin, since its encoding with smin in (1) contains no existential variables, $\{log\}$ is able to automatically compute neg(smin):
@@
   property(R,S) :- dom(R,Dr) & dom(S,Ds) & subset(Dr,Ds).where dom(R,D) encodes the domain of relation R, i.e., $dom(R)=D$. Clearly, property introduces two existential variables (Dr and Ds) thus making it impossible to use neg to compute its negation. However, the nature of these variables is different from Max in the sminsmin example. Here, these variables define the domain of R and S. In other words, Dr and Ds are sort of names for the expressions $dom(R)$ and $dom(S)$. This means that the negation of property is not:
@@
It makes no sense to state $dom(R)\neq Dr$ for all $Dr$ because it would mean that the domain of $R$ does not exist. For cases like this $\{log\}$ provides the Let predicate (Cristiá and Rossi, 2024d , Sect. 5):
@@
$\equiv\phi[e_{1}/x_{1},\dots,e_{n}/x_{n}]$ [by one-point rule (Woodcock and Davies, 1996, Sect 4.2)]
@@
$\equiv\lnot(\textsc{let~}Dr,Ds\textsc{~be~}Dr=dom(R),Ds=dom(S)\textsc{~in~}Dr\subseteq Ds)$